A restaurant in a fast food franchise has determined that the chance a customer will order a soft drink is 0.93. The probability that a customer will order a hamburger is 0.58. The probability that a customer will order french fries is 0.51. Complete parts a and b below.

a. If a customer places an​ order, what is the probability that the order will include a soft drink and no fries if these two events are​ independent?

b. The restaurant has also determined that if a customer orders a​ hamburger, the probability the customer will order fries is 0.77. Determine the probability that the order will include a hamburger and fries.
Answer #1

Let S be the event that a customer will order soft drinks.
Let H be the event that a customer will order hamburger.
Let F be the event that a customer will order french fries.

Given Information
P(S) 0.93
PH 0.58
P(F) 0.51

  • Here, we are to find:
    P(FNS
    = P(F) * PSC (Due to independence)
    =0.51 *(1-0.93
    0.0357

    Hence, the required probability is 0.0357

  • Here, it is given that,
    P(FH) = 0.77

    We are to find:
    P(FNH)
    P(FH) * PH
    = 0.77 * 0.58
    =0.4466

    Hence, the required probability is 0.4466.

I hope this clarifies your doubt. If you're satisfied with the solution, hit the Like button. For further clarification, comment below. Thank You. :)